OpenStudy (anonymous):

What is the graph of the rational function y=x^2-5x+6/x^2-1?

OpenStudy (anonymous):

You have to draw the graph?

OpenStudy (anonymous):

Yeah :/

OpenStudy (anonymous):

You need to factor the equation first

OpenStudy (anonymous):

(x-3)(x-2)/(x-1)(x+2)

OpenStudy (abmon98):

Set y=0 (x-3)(x-2)/(x+1)(x-1) =0 x-3=0 x-2=0 x=3 x=2 Set x=0 (x-3)(x-2)/(x+1)(x-1) (-3)(-2)/(1)(-1) y=-6 is the intercept Vertical asymptope what makes the denominator equal zero (x+1)(x-1)=0 x=-1 x=1 Horizontal Asymtope think about the degree x^2/x^2 both have a coeffcient of 1 1/1 =1